# V2X Cooperative Perception for Autonomous Driving: Recent Advances and Challenges

### v4

@@ V-D2 Data Privacy
Federated Learning (FL)-assisted CAVs represent an emerging paradigm aimed at mitigating privacy concerns while reducing communication costs and diversifying the seen scenarios in comparison to conventional centralized learning [276, 277]. FL enables collaborating clients to effectively train models without the need for data aggregation or direct data exchange [278]. However, determining the network resources and sensor placements for multi-stage training with multi-modal datasets in varied scenarios can be a complex challenge.
Conventional FL approaches face challenges in managing data heterogeneity arising from diverse sensor configurations. Variations in camera positions across clients can result in FL results that deviate from the optimal values of local data. To tackle this issue, the FedBEVT [251] framework introduced a federated transformer learning approach focused on camera-attentive personalization, which alleviates the negative impacts of data heterogeneity by privatizing the camera positional embedding. The work in [253] addressed the data heterogeneity in CP by implementing a novel federated dynamic weighted aggregation (FedDWA) algorithm along with a dynamic adjusting loss function. However, these methods ignore network heterogeneity. Varying connection qualities among clients can result in communication delays, thereby impeding the FL process. To fill this gap, a contextual client selection pipeline was introduced in [252], modelling networks and estimating FL communication latency under predictive transport scenarios via digital twin simulations. This FL methodology effectively tackles data and network heterogeneity, thereby improving communication efficiency through contextual client selection.
@@ V-E Perception Uncertainty
Capturing perception system uncertainties is crucial for enhancing CAV safety. The Perception Error Model (PEM) [280] incorporated single-agent perception errors into the ground truth to directly model uncertainties in the simulation pipeline, eliminating the need for synthetic sensor data. Expanding on this, the Cooperative Perception Error Model (CoPEM) [254] addresses occlusion by modelling V2X-related perception errors and uncertainties, facilitating performance tuning as sensor numbers increase. However, CoPEM requires a statistical model for each agent and overlooks the high maneuverability of vehicle targets. To resolve this, the Consensus-based Distributed Multi-Model CP (DMMCP) [255] integrated prior knowledge from diverse motion models and applied a hybrid consensus strategy with a Cubature Kalman Filter (CKF) [281] for more accurate data fusion and model uncertainty handling. Although effective, CKF initialization remains challenging, especially with uncertain initial state estimates. CMP [256] used the AB3DMOT tracker [282] to associate the detected 3D bounding boxes into trajectory segments. By incorporating cooperative perception and motion prediction aggregation, CMP is able to address the issue of propagated uncertainties and inaccuracies from the upstream detection and tracking tasks, leading to improved overall motion prediction performance.
@@
Uncertainty quantification is essential for ensuring safety in CAV systems by estimating the likelihood of outcomes when certain system aspects are uncertain. The Double-M [257] addressed uncertainty in cooperative object detection by utilizing temporal features. It applies a customized Moving-Block Bootstrap algorithm to model the multivariate Gaussian distribution for each bounding box corner, accounting for auto-correlations in time-series data. By assigning higher uncertainties to low-accuracy objects, Double-M improves accuracy, though it assumes each bounding box corner follows an independent multivariate Gaussian distribution. GevBEV [133] introduced a more comprehensive model by constructing a probabilistic BEV map using point-based spatial Gaussian distributions. It leverages Evidential Deep Learning [159] to estimate probabilities and uncertainties, interpreting Gaussian densities as support for a Dirichlet distribution. This allows GevBEV to quantify and explain prediction uncertainty, with experimental results showing it outperforms previous models on the OPV2V benchmark.
@@
In the context of object tracking, the MOT-CUP framework [258] addressed uncertainty quantification by integrating detection uncertainty into the Kalman Filter (KF) and the object association process. It models uncertainty using direct modelling [283] and conformal prediction [284], which is then applied to the Standard Deviation-based KF (SDKF) and a Negative Log-Likelihood (NLL) association improvement procedure. SDKF leverages uncertainty to enhance location prediction accuracy, while NLL improves tracking by accounting for low-quality detections. Similarly, DMSTrack [259] used a differentiable multi-sensor KF to estimate uncertainty for each detection, optimizing tracking through a learned covariance model. Nevertheless, neither approach uses multi-vehicle tracking information for improved data association, and temporary tracking failures remain a challenge. To this end, the work in [260] proposed a complementary data association module to identify and compensate for missed objects using information shared among the CAVs. Experiments show that incorporating uncertainty improves multi-object tracking performance.
@@ V-F Task Discrepancy
The spatiotemporal auto-encoder (STAR) [261] aims to disregard downstream tasks and concentrate on learning to reconstruct an entire scene through shared features in a self-supervised manner. STAR facilitates the transmission of sub-sampled features to collectively cover the entire spatial region, striking a balance between scene reconstruction performance and communication costs. The reconstructed scene can then be used for various downstream tasks without additional training. This task-agnostic methodology represents a novel direction in CP, enabling the separation of cooperation training from downstream task learning. However, due to its imperfect scene completion, STAR still exhibits a performance gap in downstream perception compared to task-specific methods.
In contrast to STAR’s self-supervised learning approach, the CORE method, as described in [262], learns to reconstruct comprehensive scenes under the guidance of BEV representations derived from the aggregation of raw sensor data from all agents. When provided with a group of agents and their 2D BEV maps, CORE conducts cooperative reconstruction through three main components: a compression module, a collaboration module, and a reconstruction module. The compression module calculates a compressed feature representation for each BEV, considering channel-wise compression and sub-sampling features along the spatial dimension. Drawing inspiration from prior work [139], CORE incorporates a lightweight attention-aware collaboration module to facilitate information aggregation. The reconstruction module employs a decoder structure to regenerate a complete scene observation from the fused features. This “learning-to-reconstruct” approach is task-agnostic and provides clear and meaningful supervision to encourage more effective cooperation, ultimately benefiting perception tasks.
@@ VI Issues for Dataset
While CP has gained momentum in recent years, it remains a nascent field. One of the key obstacles impeding the progress of CP algorithms is the substantial cost and potential safety risks associated with conducting field experiments. Such experiments necessitate deploying numerous expensive CAVs and sizeable test areas. Additionally, acquiring precise annotations requires significant labour and time, especially when dealing with CP involving multiple agents [285]. To address these challenges efficiently, an effective research strategy is to develop and validate CP algorithms within a simulated environment, which offers a cost-effective and safe alternative to real-world experimentation. However, validating V2X perception in authentic scenarios remains challenging due to the absence of well-established public benchmarks. This section delineates prominent simulation platforms and open-source datasets frequently employed in developing V2X CP systems.
@@ VI-A Simulation Tools
Numerous simulation platforms have been developed to offer multi-modal sensory inputs and 3D annotations, thereby advancing the development of multi-agent CP before the widespread availability of real-world datasets. For example, SUMO [286] is a traditional traffic and driver behaviour simulator renowned for its capacity to handle large-scale and realistic traffic flows. It incorporates dynamic modelling for each vehicle, enabling users to swiftly create customized traffic scenarios through the Traffic Control Interface (TraCI) API. SUMO is highly convenient for integrating naturalistic trajectory data, which can be directly applied to the surrounding environment for testing CDA. However, it has limitations when providing high-resolution sensor data with exceptional fidelity.
@@ VI-A2 OpenCDA
OpenCDA [288, 289] stands as an open and dedicated co-simulation framework designed for CDA. It encompasses both traffic and vehicles, offering a complete CDA vehicle software pipeline. OpenCDA seamlessly integrates CARLA and SUMO for lifelike scene rendering, vehicle modelling, and traffic simulation. This framework facilitates the evaluation of CDA within a CARLA + SUMO co-simulation environment and supports diverse levels and categories of cooperation among CAVs during simulation. OpenCDA also provides a comprehensive library of CDA research pipelines, source code, and standard autonomous driving components encompassing perception, localization, planning, control, and V2X communication modules. One of its key advantages is its flexibility, allowing users to easily substitute default modules with custom-designed ones to assess the impact of new functionalities on overall CDA performance.
@@ VI-A5 CMM Co-simulation Platform
TABLE VII: Comparison between Cooperative Perception-related Datasets.
Category Dataset Year Scenario Agent Number Annotation Range Object Category RGB Image LiDAR Frame 3D Box Maps OD OT SS MP AP
@@
V2X-Seq [293] 2023 V2I 2 280m 10 15k 15k -- $\surd$ $\surd$ $\surd$ $\times$ $\surd$ $\times$
@@
LUCOOP [296] 2023 V2V 3 200m 5 0 54k -- $\surd$ $\surd$ $\surd$ $\times$ $\times$ $\times$
@@
RCooper [298] 2024 I2I 2 230m 10 50k 30k -- $\times$ $\surd$ $\surd$ $\times$ $\times$ $\times$
@@ VI-B Open Source Dataset
The development of CP models relies on perception datasets gathered from various onboard sensors like radar, camera, and LiDAR. Utilizing these sensors for real-world perception tasks can be costly and time-consuming. A more economically viable alternative is to obtain cooperative data from simulation platforms featuring high-fidelity sensor modelling and perception capabilities. This subsection will outline several datasets relevant to CDA research, summarized in Table VII.
@@ VI-B2 Simulation Dataset
While realistic datasets allow CP models to generalize effectively to real-world driving situations, they are expensive, time-consuming, and limited in scenarios. A cost-effective alternative is to gather large-scale datasets in a high-fidelity simulator. The Cooperative Driving Dataset (CODD) [301], generated using the CARLA simulator, comprises 108 snippets, each with 125 temporal frames of LiDAR point cloud data. Each frame contains LiDAR data from all vehicles in the scenario, sensor poses, and ground-truth annotations for 3D bounding boxes of cars and pedestrians. CODD focuses on a single driving scenario. In contrast, the OPV2V dataset [302] includes both camera images and LiDAR point clouds from connected vehicles across over 70 diverse scenes, spanning eight CARLA towns and Culver City, Los Angeles, using OpenCDA and CARLA. Both CODD and OPV2V provide ground truth for V2V cooperative 3D object detection, but neither includes roadside infrastructure, limiting their ability to evaluate V2I perception [308].
@@ VII-A1 Detection Performance Comparison
To ensure fairness and generalizability, all reported results reflect the best outcomes from experimental evaluations of existing methods, with only those methods assessed on at least two datasets included. Table VIII displays the AP values at IoU thresholds of 0.5 and 0.7 across five datasets under default settings, highlighting the highest AP values and indicating the source for each result. The observed performance differences across methods and datasets indicate that certain algorithms are more effective under specific environmental or data conditions.
Table VIII results indicate that DI-V2X achieves the highest AP@0.5 on DAIR-V2X at 0.7882, while DiscoNet leads on V2V4Real with an AP@0.5 of 0.7360. For AP@0.7 on V2V4Real, both FPV-RCNN and MACP achieve top performance at 0.4790, though MACP scores slightly lower than FPV-RCNN with an AP@0.5 of 0.6760. Among simulation-based datasets, MKD-Cooper performs exceptionally on V2X-Sim, with an AP@0.5 of 0.9300 and AP@0.7 of 0.8520, and achieves a high AP@0.7 of 0.9240 on OPV2V, positioning it as a leading method in simulated environments. V2X-PC demonstrates robust performance across both real and simulated datasets, achieving an AP@0.5 of 0.7689 and AP@0.7 of 0.6939 on DAIR-V2X and even higher scores on V2XSet with an AP@0.5 of 0.9283 and AP@0.7 of 0.8955. These findings suggest that V2X-PC excels across diverse datasets, indicating strong adaptability to various conditions.
TABLE VIII: 3D Object Detection Performance Comparison On DAIR-V2X, V2V4Real, OPV2V, V2XSet, and V2X-Sim datasets, with AP@0.5 and AP@0.7.
@@
AP@0.5/0.70.50.70.5/0.70.5 / 0.7 Source AP@0.5/0.70.50.70.5/0.70.5 / 0.7 Source AP@0.5/0.70.50.70.5/0.70.5 / 0.7 Source AP@0.5/0.70.50.70.5/0.70.5 / 0.7 Source AP@0.5/0.70.50.70.5/0.70.5 / 0.7 Source
@@
CoAlign [190] 2023 $0.7460/0.6040$ [190] $0.7090/0.4170$ [193] 0.9660$/0.9120$ [190] $0.9190/0.8050$ [191] $0.8580/0.7650$ [190]
V2VFormer [216] 2024 - - 0.6540///0.3610 [216] $0.9170/0.8760$ [216] - - 0.6150///0.5500 [216]
CoBEVGlue [193] 2024 0.7400///0.5820 [193] 0.7020///0.4310 [193] 0.9580///0.9090 [193] - - - -
IFTR [134] 2024 0.2058///0.0665 [134] - - 0.8556///0.6604 [134] 0.7173///0.4967 [134] - -
V2X-M2C [232] 2024 - - - - 0.8870///0.8280 [232] 0.9220///0.8600 [232] - -
CPSC [142] 2024 0.6542///0.4784 [142] - - 0.8911///0.7797 [142] - - - -
V2X-INCOP [236] 2024 0.6548///0.5526 [236] - - 0.9256///0.8318 [236] - - - -
RoCo[191] 2024 0.7630///0.6200 [191] - - - - 0.9190///0.8050 [191] - -
MACP [228] 2024 - - 0.6760///0.4790 [228] 0.9370/0.9030 [228] - - - -
DI-V2X [229] 2024 0.7882///0.6616 [229] - - - - 0.9270///0.8270 [229] - -
MKD-Cooper [218] 2024 - - - - 0.9650/0.9650/0.9650 /0.9240 [218] - - 0.9300///0.8520 [218]
V2X-PC [136] 2024 0.7689///0.6939 [136] - - - - 0.9283///0.8955 [136] - -

### v5

@@ V-D2 Data Privacy
Federated Learning (FL)-assisted CAVs represent an emerging paradigm aimed at mitigating privacy concerns while reducing communication costs and diversifying the seen scenarios in comparison to conventional centralized learning [182]. FL enables collaborating clients to effectively train models without the need for data aggregation or direct data exchange [183]. However, determining the network resources and sensor placements for multi-stage training with multi-modal datasets in varied scenarios can be a complex challenge.
Conventional FL approaches have been challenged by data heterogeneity resulting from diverse sensor configurations. Variations in camera positions across clients have caused FL outcomes to deviate from the optimal values associated with local data. To address this issue, a federated transformer learning framework, FedBEVT [164], was introduced, in which camera-attentive personalization was applied to alleviate the negative effects of data heterogeneity by privatizing the camera positional embedding. In [166], a Federated Dynamic Weighted Aggregation (FedDWA) algorithm, along with a dynamically adjusted loss function, was implemented to manage data heterogeneity in CP. However, network heterogeneity was not considered in these methods. Communication delays caused by varying connection qualities among clients have hindered the efficiency of the FL process. To overcome this limitation, a contextual client selection pipeline was proposed in [165], where network conditions were modeled and FL communication latency was estimated under predictive transport scenarios using digital twin simulations. Through this approach, both data and network heterogeneity were effectively addressed, and communication efficiency was enhanced via context-aware client selection.
@@ V-E Perception Uncertainty
Capturing perception system uncertainties is crucial for enhancing CAV safety. The Perception Error Model (PEM) [185] was developed to incorporate single-agent perception errors into the ground truth, thereby enabling the direct modeling of uncertainties within the simulation pipeline and eliminating the reliance on synthetic sensor data. This approach was further extended in the Cooperative Perception Error Model (CoPEM) [167], which accounted for occlusion-related perception errors in vehicle-to-everything communication scenarios. CoPEM enabled performance tuning as the number of sensors increased. However, it required a statistical model for each agent and did not consider the high maneuverability of vehicle targets. To address this limitation, the Consensus-based Distributed Multi-Model Cooperative Perception (DMMCP) framework [168] was proposed. It integrated prior knowledge from multiple motion models and adopted a hybrid consensus strategy using the Cubature Kalman Filter (CKF) [186] to improve data fusion accuracy and manage model uncertainty. Despite its effectiveness, the initialization of the CKF posed challenges, particularly when initial state estimates were uncertain. More recently, CMP [169] was introduced, which associated detected three-dimensional bounding boxes with trajectory segments. By aggregating cooperative perception and motion prediction data, CMP effectively mitigated propagated uncertainties and inaccuracies from upstream detection and tracking stages, resulting in enhanced motion prediction performance.
@@
Uncertainty quantification is regarded as essential for ensuring safety in connected and autonomous vehicle (CAV) systems, as it allows the likelihood of outcomes to be estimated when aspects of the system remain uncertain. In Double-M [170], temporal features were utilized to address uncertainty in cooperative object detection. A customized Moving-Block Bootstrap algorithm was employed to model the multivariate Gaussian distribution for each bounding box corner, with auto-correlations in time-series data taken into account. Higher uncertainties were assigned to low-accuracy objects, resulting in improved detection accuracy. However, it was assumed that each bounding box corner followed an independent multivariate Gaussian distribution. A more comprehensive uncertainty modeling approach was proposed in GevBEV [47], in which a probabilistic BEV map was constructed using point-based spatial Gaussian distributions. Evidential Deep Learning [70] was applied to estimate predictive probabilities and associated uncertainties, where the Gaussian densities were interpreted as evidence supporting a Dirichlet distribution. This allowed prediction uncertainty to be quantified and explained, and superior performance was demonstrated on the OPV2V benchmark.
@@
In the context of object tracking, uncertainty quantification was addressed in the MOT-CUP framework [171] by integrating detection uncertainty into the Kalman Filter (KF) and the object association process. Uncertainty was modeled using direct modeling techniques [187] and conformal prediction methods [188], and subsequently applied to the Standard Deviation-based KF (SDKF) and a Negative Log-Likelihood (NLL) association refinement procedure. In SDKF, uncertainty was leveraged to enhance the accuracy of location prediction, while in NLL, tracking performance was improved by accounting for low-quality detections. Similarly, in DMSTrack [172], a differentiable multi-sensor KF was utilized to estimate uncertainty for each detection, with a learned covariance model used to optimize tracking. However, multi-vehicle tracking information was not incorporated in either approach to enhance data association, and temporary tracking failures remained unresolved. To address this, a complementary data association module was introduced in [173] to identify and recover missed objects using information shared among CAVs. Experimental results demonstrated that the incorporation of uncertainty improved multi-object tracking performance.
@@ V-F Task Discrepancy
The Spatiotemporal Auto-Encoder (STAR) [174] was designed to disregard downstream tasks and focus on reconstructing the entire scene through shared features in a self-supervised manner. In this framework, sub-sampled features are transmitted among agents to collectively cover the complete spatial region, allowing a balance to be maintained between scene reconstruction performance and communication costs. The reconstructed scene can subsequently be utilized for various downstream tasks without requiring additional training. This task-agnostic approach has been proposed as a novel direction in CP, enabling the decoupling of cooperation training from downstream task learning. However, due to incomplete scene reconstruction, a performance gap is still observed in downstream perception when compared to task-specific methods.
In contrast to STAR’s self-supervised learning approach, the CORE method [175], learns to reconstruct comprehensive scenes under the guidance of BEV representations derived from the aggregation of raw sensor data from all agents. When provided with a group of agents and their 2D BEV maps, CORE conducts cooperative reconstruction through three main components: a compression module, a collaboration module, and a reconstruction module. The compression module calculates a compressed feature representation for each BEV, considering channel-wise compression and sub-sampling features along the spatial dimension. CORE incorporates a lightweight attention-aware collaboration module to facilitate information aggregation. The reconstruction module employs a decoder structure to regenerate a complete scene observation from the fused features. This “learning-to-reconstruct” approach is task-agnostic and provides clear and meaningful supervision to encourage more effective cooperation, ultimately benefiting perception tasks.
@@ VI Issues for Dataset
While CP has gained momentum in recent years, it remains a nascent field. One of the key obstacles impeding the progress of CP algorithms is the substantial cost and potential safety risks associated with conducting field experiments. Such experiments necessitate deploying numerous expensive CAVs and sizeable test areas. Additionally, acquiring precise annotations requires significant labour and time, especially when dealing with CP involving multiple agents. To address these challenges efficiently, an effective research strategy is to develop and validate CP algorithms within a simulated environment, which offers a cost-effective and safe alternative to real-world experimentation. However, validating V2X perception in authentic scenarios remains challenging due to the absence of well-established public benchmarks. This section delineates prominent simulation platforms and open-source datasets frequently employed in developing V2X CP systems.
@@ VI-A Simulation Tools
Numerous simulation platforms have been developed to offer multi-modal sensory inputs and 3D annotations, thereby advancing the development of multi-agent CP before the widespread availability of real-world datasets. For example, Simulation of Urban MObility (SUMO) [189] is a traditional traffic and driver behaviour simulator renowned for its capacity to handle large-scale and realistic traffic flows. It incorporates dynamic modelling for each vehicle, enabling users to swiftly create customized traffic scenarios through the Traffic Control Interface (TraCI) API. SUMO is highly convenient for integrating naturalistic trajectory data, which can be directly applied to the surrounding environment for testing CDA. However, it has limitations when providing high-resolution sensor data with exceptional fidelity.
@@ VI-A2 OpenCDA
OpenCDA [191] stands as an open and dedicated co-simulation framework designed for CDA. It encompasses both traffic and vehicles, offering a complete CDA vehicle software pipeline. OpenCDA seamlessly integrates CARLA and SUMO for lifelike scene rendering, vehicle modelling, and traffic simulation. This framework facilitates the evaluation of CDA within a CARLA + SUMO co-simulation environment and supports diverse levels and categories of cooperation among CAVs during simulation. OpenCDA also provides a comprehensive library of CDA research pipelines, source code, and standard autonomous driving components encompassing perception, localization, planning, control, and V2X communication modules. One of its key advantages is its flexibility, allowing users to easily substitute default modules with custom-designed ones to assess the impact of new functionalities on overall CDA performance.
@@ VI-A5 CMM Co-simulation Platform
TABLE VI: Comparison between Cooperative Perception-related Datasets.
Category Dataset Year Scenario
Agent
Number
Annotation
Range
Object
Category
RGB
Image
LiDAR
Frame
3D
Box
Maps
OD
OT
$S$
MP
AP
@@
V2X-Seq [195] 2023 V2I 2 280m 10 15k 15k - $\surd$ $\surd$ $\surd$ $\times$ $\surd$ $\times$
@@
LUCOOP [198] 2023 V2V 3 200m 5 0 54k - $\surd$ $\surd$ $\surd$ $\times$ $\times$ $\times$
@@
RCooper [200] 2024 I2I 2 230m 10 50k 30k - $\times$ $\surd$ $\surd$ $\times$ $\times$ $\times$
@@ VI-B Open Source Dataset
The development of CP models relies on perception datasets gathered from various onboard sensors like radar, camera, and LiDAR. Utilizing these sensors for real-world perception tasks can be costly and time-consuming. A more economically viable alternative is to obtain cooperative data from simulation platforms featuring high-fidelity sensor modelling and perception capabilities. This subsection will outline several datasets relevant to CDA research, summarized in Table VI.
@@ VI-B2 Simulation Dataset
While realistic datasets allow CP models to generalize effectively to real-world driving situations, they are expensive, time-consuming, and limited in scenarios. A cost-effective alternative is to gather large-scale datasets in a high-fidelity simulator. The Cooperative Driving Dataset (CODD) [203], generated using the CARLA simulator, comprises 108 snippets, each with 125 temporal frames of LiDAR point cloud data. Each frame contains LiDAR data from all vehicles in the scenario, sensor poses, and ground-truth annotations for 3D bounding boxes of cars and pedestrians. CODD focuses on a single driving scenario. In contrast, the OPV2V dataset [204] includes both camera images and LiDAR point clouds from connected vehicles across over 70 diverse scenes, spanning eight CARLA towns and Culver City, Los Angeles, using OpenCDA and CARLA. Both CODD and OPV2V provide ground truth for V2V cooperative 3D object detection, but neither includes roadside infrastructure, limiting their ability to evaluate V2I perception.
@@ VII-A1 Detection Performance Comparison
To ensure fairness and generalizability, all reported results reflect the best outcomes from experimental evaluations of existing methods, with only those methods assessed on at least two datasets included. Table VII displays the AP values at IoU thresholds of 0.5 and 0.7 across five datasets under default settings, highlighting the highest AP values and indicating the source for each result. The observed performance differences across methods and datasets indicate that certain algorithms are more effective under specific environmental or data conditions.
Table VII results indicate that DI-V2X achieves the highest AP@0.5 on DAIR-V2X at 0.7882, while DiscoNet leads on V2V4Real with an AP@0.5 of 0.7360. For AP@0.7 on V2V4Real, both FPV-RCNN and MACP achieve top performance at 0.4790, though MACP scores slightly lower than FPV-RCNN with an AP@0.5 of 0.6760. Among simulation-based datasets, MKD-Cooper performs exceptionally on V2X-Sim, with an AP@0.5 of 0.9300 and AP@0.7 of 0.8520, and achieves a high AP@0.7 of 0.9240 on OPV2V, positioning it as a leading method in simulated environments. V2X-PC demonstrates robust performance across both real and simulated datasets, achieving an AP@0.5 of 0.7689 and AP@0.7 of 0.6939 on DAIR-V2X and even higher scores on V2XSet with an AP@0.5 of 0.9283 and AP@0.7 of 0.8955. These findings suggest that V2X-PC excels across diverse datasets, indicating strong adaptability to various conditions.
TABLE VII: 3D Object Detection Performance Comparison On DAIR-V2X, V2V4Real, OPV2V, V2XSet, and V2X-Sim datasets, with AP@0.5 and AP@0.7.
@@
AP@$0.5/0.7$ Source AP@$0.5/0.7$ Source AP@$0.5/0.7$ Source AP@$0.5/0.7$ Source AP@$0.5/0.7$ Source
@@
CoAlign [102] 2023 $0.7460/0.6040$ [102] $0.7090/0.4170$ [105] 0.9660/0.9120/0.9120 [102] $0.9190/0.8050$ [103] $0.8580/0.7650$ [102]
V2VFormer [129] 2024 - - 0.6540//0.3610 [129] $0.9170/0.8760$ [129] - - 0.6150//0.5500 [129]
CoBEVGlue [105] 2024 0.7400//0.5820 [105] 0.7020//0.4310 [105] 0.9580//0.9090 [105] - - - -
IFTR [48] 2024 0.2058//0.0665 [48] - - 0.8556//0.6604 [48] 0.7173//0.4967 [48] - -
V2X-M2C [145] 2024 - - - - 0.8870//0.8280 [145] 0.9220//0.8600 [145] - -
CPSC [56] 2024 0.6542//0.4784 [56] - - 0.8911//0.7797 [56] - - - -
V2X-INCOP [149] 2024 0.6548//0.5526 [149] - - 0.9256//0.8318 [149] - - - -
RoCo[103] 2024 0.7630//0.6200 [103] - - - - 0.9190//0.8050 [103] - -
MACP [141] 2024 - - 0.6760//0.4790 [141] 0.9370/0.9030 [141] - - - -
DI-V2X [142] 2024 0.7882//0.6616 [142] - - - - 0.9270//0.8270 [142] - -
MKD-Cooper [131] 2024 - - - - 0.9650/0.9650/0.9240 [131] - - 0.9300//0.8520 [131]
V2X-PC [50] 2024 0.7689//0.6939 [50] - - - - 0.9283//0.8955 [50] - -